Grade 11Mechanics

a balloon is moving with a velocity of 5m/s in upward direction. at t=0 the balloon is at the ground. at t= 10s a ball is left gently from the balloon . the speed of the ball at the ground is

annabeth
8 Years agoGrade 11

1 Answer

NIKHIL GHODKE
8 Years ago
In this case the balloon is moving with constant velocity of 5m/s sec in upward direction
so as it is moving with constant velocity we do not consider the effect of gravity on balloon.
At 10 sec the the distance covered by balloon =u*t
= 5*10
= 50m
now as the ball was moving with the balloon ,so the ball even acquires the vertical velocity of the balloon which is 5m/s upward.
now using the equation
v2=u2 +2as
u=+5m/s
a=g=–9.8m/s2
s= -50m(take not eof the negative sign  in this sum we have taken vectors away from ground positive and towards ground negative ).
v2= 25+2*(-9.8)*(-50)
v2 = 1005
v = (1005)0.5
I TRULY ENCOURAGE TO U TO ASK SUCH DOUBT AS  no doubt is silly or small
comment if have any problem